OpenStudy (anonymous):

simplify the expression square root of 27 times square root of 2

OpenStudy (anonymous):

\[\sqrt{27}\times \sqrt{2} \]

jhonyy9 (jhonyy9):

3sqrt3 *sqrt2 = 3sqrt6

hero (hero):

= sqrt{27*2} = sqrt{54} = sqrt{9*6} = sqrt{9} * sqrt{6} = 3 sqrt{6}

OpenStudy (anonymous):

how did you get 3 square roots of three

jhonyy9 (jhonyy9):

27=9*3 sqrt(9*3)=3sqrt3

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

so take the perfact square of 9 out of 27 so 9 times 3 =27 then the suare root of 9 is 3 so 3 suare roots of the 3 that cant be simplified right

jhonyy9 (jhonyy9):

yes

hero (hero):

If you use my approach, you wouldn't have to worry about running into 3 sqrt{3}

OpenStudy (anonymous):

ok ya i think Hero's way makes a little more sence. thanks though

OpenStudy (anonymous):

so 27 times 2 =54 than take the perfect square root of 9 out so 9 times 6= 54 then simplify so 3 square roots of 6 corect?

OpenStudy (anonymous):

so as another example \[\sqrt{10} \times \sqrt{8}\] 10 times 8 is 80 then take 16 out so 16 times 5 = 80 then simplify to 4 square roots of 5 is that how it works?

hero (hero):

yup

hero (hero):

good job

OpenStudy (anonymous):

ok but what about \[\sqrt{12} \times \sqrt{18}\] i have a fealing theres an extra step i don't know how to do

hero (hero):

It's the same strategy...look for the highest square, then split it up into a square root of two numbers, one being a square.

hero (hero):

If you multiply 12*18 you get 216. 216/6 = 36 so do sqrt{36*6}

OpenStudy (anonymous):

i got 18 times 12 is 216 then 9 times 24 so 3 is from the 9 and the 24 also can go to 4times 6 which is 2 square roots of 6 so would i add 3+ 2

OpenStudy (anonymous):

and make it 5square roots of 6

hero (hero):

Basically, there are two strategies to finding the square. You either divide by multiples of two until you find a square, or you divide by squares until you get a non-decimal integer.

hero (hero):

Basically, I did this: 216/2 no 216/4 no 216/6 yes Or you can find the nearest square and divide by that like so: 216/196 no 216/169 no ... 216/36 yes It depends on which number you're working with but either way works.